OpenStudy (anonymous):

A polygon has an area of 144 square inches and one of its sides is 8 inches long. If a second similar polygon has an area of 64 square inches, what is the length of the corresponding side in the second polygon? 5 1/3

OpenStudy (anonymous):

so im looking for number that will equal 512

OpenStudy (anonymous):

i get 3.5

OpenStudy (anonymous):

wel 3.6 isnt an option

OpenStudy (anonymous):

wel back to the top thank you

Directrix (directrix):

I got 16/3 or 5.3. Maybe I messed up. I'll post my work.

OpenStudy (anonymous):

yeah no

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

6 5/9 is another option and 12 and 18 which i dont even look at so either of of those two fractions

Directrix (directrix):

If two polygons are similar, the square of the scale factor of the two polygons is equal to the ratio of the areas of the polygon. That said, the square of the ratio of x to 8 is equal to the ratio of the areas of the similar polygons. 64/144 = (x/8)^2 8/12 = x/8 2/3 = x/8 3x = 16 x = 16/3 = 5 1/3

OpenStudy (anonymous):

that means im not that off thank you

OpenStudy (mertsj):

I overlooked that area part. I was thinking perimeter.

OpenStudy (mertsj):

Sorry